(solve) > Implement the boolean function F(x,y,z) = xy + x'y + yz

write all the steps
and identify the law rule
Answer #1
Answer:
F(x,y,z) = xy + x'y + yz
Applying distributive law
F(x,y,z) = (x+x')y + yz
Applying inverse law
F(x,y,z) = (1)y + yz
Applying identity law
F(x,y,z) = y + yz
Applying distributive law
F(x,y,z) = y(1 + z)
Applying identity law
F(x,y,z) = y(1)
Applying identity law
F(x,y,z) = y